OpenStudy (anonymous):

graph the parabola y=x^2+3

jagr2713 (jagr2713):

Hey you can use a graphing calculator here https://www.desmos.com/calculator

OpenStudy (anonymous):

I did it and it only showed me one point and not the other points

jagr2713 (jagr2713):

Ok put in the equation in the box

OpenStudy (anonymous):

boom bo bo

OpenStudy (anonymous):

b

OpenStudy (anonymous):

b

jagr2713 (jagr2713):

@henryl Please dont spam this post.

jagr2713 (jagr2713):

@atlantaslast did u put the equation in there

OpenStudy (anonymous):

I did, I typed in y=x^2+3 and it only showed me one point which is 0,3

OpenStudy (anonymous):

okay

jagr2713 (jagr2713):

Here let me make a vid

OpenStudy (anonymous):

i will not spam the post

jagr2713 (jagr2713):

Thanks

jagr2713 (jagr2713):

Here @atlantaslast http://gyazo.com/18f99dd1bec6c5aac18a429d822f1baf

OpenStudy (anonymous):

Thank you so much :)

OpenStudy (anonymous):

\[\frac{ 56 }{ 92 }=\lfloor 56\left| 490\int\limits_{123}^{1034\prod_{67}^{34}} \right| \rfloor\]

jagr2713 (jagr2713):

Anytime :D

OpenStudy (anonymous):

prod

jagr2713 (jagr2713):

@henryl this is your second and final warning

OpenStudy (anonymous):

\[\prod_{}^{}\]

OpenStudy (anonymous):

srry ddidnt see that